# OMP4Py: a pure Python implementation of OpenMP

### v1

@@ 4.1 Numerical algorithms
All the source codes can be found in the OMP4Py repository. Figure 8 shows the execution times of the different parallel benchmarks using from 1 to 24 threads. It can be observed that the scalability of the OMP4Py applications is not good. For instance, only 4 out of 6 benchmarks benefit from using 4 threads, while using more threads consistently proves detrimental to performance. The best speedup overall is 3.18$\times$, obtained for the molecular dynamics application and 4 threads. It is important to highlight that these scalability problems are not related to the Python code generated by OMP4Py. As we will show in Section 4.2, non-numerical OMP4Py applications exhibit strong scalability with codes containing the same structure and OpenMP directives as those discussed in this section. Therefore, these results demonstrate that the current version of the Python interpreter (v3.13) still lacks mature support for multithreading, with several open implementation issues that need to be addressed to exploit it efficiently. While this new version is an important step forward in removing the GIL, it still limits the use of OMP4Py to a small number of threads when running numerical applications.
@@
Finally, for illustrative purposes, Figure 8 also shows the performance results obtained by the applications when compiled using PyOMP (Numba). As expected, they are significantly faster than pure Python, as Numba compiles Python code into optimized machine code, bypassing the slower interpreted execution of standard Python. This leads to substantial reductions in the execution times, particularly for compute-intensive tasks such as numerical computations. In any case, we have observed problems in the scalability of some benchmarks (see fft and md results). Additionally, PyOMP was unable to execute the jacobi method using more than one thread.
@@ 4.2 Applications using non-numerical libraries and Python objects
Figure 9 shows the scalability of the Graph Clustering and Wordcount applications using up to 48 threads. The scalability of OMP4Py is good, especially for Wordcount, which achieves a 25.5$\times$ speedup compared to the sequential execution when using 48 threads. This behavior is very different from the performance results obtained for the numerical applications (see Figure 8). However, the structure and OpenMP directives used in both types of applications are very similar, with the only differences being related to the operations performed inside the loops. For example, when closely examining the OMP4Py codes in Figure 10 for the $\pi$ calculation (left) and the Wordcount application (right), it can be observed that the OpenMP clauses are identical, with only the operations in the $i$ loop and the critical block differing. As a result, the scalability issues detected for the $\pi$ calculation cannot be attributed to OMP4Py, since the code generated for both applications is essentially the same. Therefore, these results can only be explained by the fact that free-threaded support for Python 3.13 is still in an experimental stage, leaving plenty of room for improving performance and efficiency.
1from omp4py import *
2
3@omp
4def pi(n):
5 w = 1.0 / n
6 PI = 0.0
7 with omp("parallel"):
8 pi_local = 0.0
9 with omp("for"):
10 for i in range(n):
11 local = (i + 0.5) * w
12 pi_local += 4.0 / (1.0 + local * local)
13 with omp("critical"):
14 PI += pi_local
15
16 return PI * w
@@ 14 PI += pi_local
Figure 10: OMP4Py user codes for the Pi (left) and Wordcount (right) applications.
@@ 5 Conclusions
In this paper, we introduce OMP4Py [^4] It is publicly available at https://github.com/citiususc/omp4py, a novel implementation of the OpenMP standard designed specifically for Python. It fully supports the API specifications of version 3.0. OMP4Py integrates OpenMP into Python by adapting its directive-based approach, allowing users to embed parallel constructs directly into their code. These transformer directives enable parallel execution by instructing OMP4Py to modify the Python code as needed.
The performance evaluation of OMP4Py reveals limited scalability when using numerical algorithms. Specifically, only 4 out of 6 benchmarks showed improvement with 4 threads, with performance typically degrading when more threads were used. However, we demonstrated that these scalability issues are due to Python’s multithreading limitations in version 3.13, rather than OMP4Py’s implementation.
@@
In the future, OMP4Py will be extended to support newer versions of the OpenMP standard, including versions 4.0 through 6.0. This will involve incorporating advanced features such as task dependencies and thread teams. Additionally, the current implementation will be optimized by reducing mutex locks and exploring atomic operations, which are commonly used in C-based OpenMP implementations.

### v2

@@ 4.1 Numerical algorithms
All the source codes can be found in the OMP4Py repository. Figure 8 shows the execution times of the different parallel benchmarks using from 1 to 24 threads. It can be observed that the scalability of the OMP4Py applications is not good. For instance, only 4 out of 6 benchmarks benefit from using 4 threads, while using more threads consistently proves detrimental to performance. The best overall speedup, calculated as the ratio of sequential time to parallel time, is 3.18$\times$, obtained for the molecular dynamics application using 4 threads.
To better understand the behavior of OMP4Py parallel applications, as an illustrative example, we will focus on tracking core usage while executing the pi application with different thread counts (Figure 9). The profiling was performed using the mpstat command, which is part of the sysstat package. This tool provides detailed statistics on CPU usage, allowing us to monitor individual core usage and overall system performance. The data was captured at 5-second intervals. As observed, only with 1, 2, and 4 threads does the average core usage reach 100% or close to it, which corresponds to the cases where the parallel code scales efficiently (see Figure 8). However, a significant decrease in core usage is detected as the number of threads increases. Specifically, average values of only about 54% and 22% were obtained when using 8 and 16 threads, respectively. The fact that the average core usage remains low suggests that thread synchronization overhead (e.g., waiting for shared resources) likely prevents the parallel application from scaling effectively. Note that this behavior was already identified by Python developers and currently remains an open issue  [^4] https://github.com/python/cpython/issues/118649 [online, accessed February 27, 2025]. In any case, as we will demonstrate next through several experiments, it is not related to the Python code generated by OMP4Py.
Figure 9: Average core usage when running the pi application using different number of threads. Time step is 5 seconds.
First, we aim to estimate the baseline overhead introduced by OMP4Py. To do so, we measured the execution times of numerical applications running with OMP4Py using a single thread and in a purely sequential manner, without any explicit reference to OMP4Py in the user code. Ten measurements were carried out for each version and application. The execution time differences between the OMP4Py (single-thread) and sequential implementations vary slightly across applications but remain consistently low, with all differences below 0.2%. The results of the Wilcoxon rank-sum test indicate that these differences are not statistically significant in all cases, meaning OMP4Py does not introduce a meaningful impact on execution time.
Once we established that the overhead introduced by OMP4Py is negligible, we will investigate the origin of the poor scalability by implementing, as a representative case, three different versions of the pi application. The user codes are shown in Figure 10. The first approach (top code) is a version in which thread creation, loop iteration distribution across threads, and final reduction of the $\pi$ value are all handled automatically by OMP4Py. The second approach (middle code) replaces the automatic distribution of loop iterations performed by OMP4Py with manual distribution. While thread creation and the final reduction are still managed by OMP4Py, the code explicitly calculates the start and end indices for each thread. The compute function is called (lines 1–5, middle code), taking the calculated range and iterating through it to perform the $\pi$ calculation. The workload division is based on the number of threads, k, and each thread’s identifier, tid, which are retrieved using the omp4py.omp_get_num_threads() and omp4py.omp_get_thread_num() functions (lines 12–13, middle code). This approach represents the explicit version of the code that OMP4Py would have generated automatically in the first approach. The third version (bottom code) builds upon the second version but incorporates Cython [15] to optimize the compute routine. Cython compiles Python code into highly optimized C code, which is then compiled into a shared library. This allows Python to call the function as a native extension, bypassing the Python interpreter and significantly improving performance. By applying the @cython.compile decorator (line 1, bottom code), the computation of $\pi$ is further accelerated, as the compute function is compiled into efficient C code. Note that thread creation and reduction for computing $\pi$ remain automatically managed by OMP4Py, while loop iteration distribution remains manual. This is because the omp4py_pi function is identical to the second version but with Cython’s performance enhancements.
1@omp
2def omp4py_pi(n):
3 w = 1.0 / n
4 PI = 0.0
5 with omp("parallel for reduction(+:PI)"):
6 for i in range(n):
7 local = (i + 0.5) * w
8 PI += 4.0 / (1.0 + local * local)
9 return PI * w
1def compute(start, end, w, PI):
2 for i in range(start, end):
3 local = (i + 0.5) * w
4 PI += 4.0 / (1.0 + local * local)
5 return PI
6
7@omp
8def omp4py_pi(n):
9 w = 1.0 / n
10 PI = 0.0
11 with omp("parallel reduction(+:PI)"):
12 k = omp4py.omp_get_num_threads()
13 tid = omp4py.omp_get_thread_num()
14 chunk = n // k
15 rem = n % k
16
17 start = chunk * tid
18 end = start + chunk + (1 if tid < rem else 0)
19 PI = compute(start, end, w, PI)
20 return PI * w
1@cython.compile
2def compute(start: cython.long, end: cython.long,
3 w:cython.float, PI:cython.float) -> cython.float:
4 i: cython.long
5 for i in range(start, end):
6 local:cython.float = (i + 0.5) * w
7 PI += 4.0 / (1.0 + local * local)
8 return PI
Figure 10: User codes for the pi application: (top) OMP4Py (middle) manually parallelized, and (bottom) compute function using Cython.
Figure 11 presents the execution times of the three code versions of the pi benchmark, using 1 to 24 threads. Note that the results for the OMP4Py version (top code in Figure 10) are the same as those displayed in Figure 8 for the pi application. From these results, several conclusions can be drawn:
  • 1.
    Since the performance of the OMP4Py version and the manually parallelized version are identical, it can be concluded that the overhead of the automatic loop iteration distribution performed by OMP4Py is negligible. Furthermore, scalability issues persist in both versions, indicating that they are not caused by work distribution.
  • 2.
    As expected, the execution times when using Cython to compile the compute function are significantly faster than when using pure Python. We must highlight that, with one thread, pure Python takes about 785 seconds to calculate $\pi$, whereas the Cython version requires only about 10 seconds. Since the pi application is a compute-intensive benchmark, this behavior confirms that these performance differences are solely due to the fundamental differences between compiled (C) and interpreted (Python) languages.
  • 3.
    The scalability of the Cython code is excellent; there are no signs of scalability issues. For example, the speedup reaches 21.7$\times$ when using 24 threads. Considering that the only difference between the manually parallelized version (middle code in Figure 10) and the Cython version is the compute routine, we can conclude that the scalability problems are not related to OMP4Py, since the overhead of thread creation, loop iteration distribution, and final reduction remains the same. Therefore, the only plausible explanation is that inefficiencies exist in the Python interpreter when multiple threads execute the compute routine concurrently. In any case, these issues cannot be attributed to the code generated by OMP4Py.
Figure 11: Scalability of the different user codes shown in Figure 10 for the pi application. Axis in log scale.
The previous analysis demonstrates that the current version of the Python interpreter (v3.13) still lacks mature support for multithreading, with several unresolved implementation issues that hinder its efficient use. In particular, as of February 2025, Python developers are dealing with over 100 open issues labeled under the free-threading topic—half of which are bugs, and 15 of which correspond to critical crashes. While this new version is an important step forward in removing the GIL, it still limits the use of OMP4Py to a small number of threads when running numerical (compute-intensive) applications. However, it is important to highlight that as these issues in the Python interpreter are progressively resolved, the scalability limitations of OMP4Py when running numerical applications will gradually disappear. On the other hand, as we will show in Section 4.2, non-numerical OMP4Py applications exhibit strong scalability with codes containing the same structure and OpenMP directives as those discussed in this section.
@@ 19 PI = compute(start, end, w, PI)
Finally, for illustrative purposes, Figure 8 also shows the performance results obtained by the applications when compiled using PyOMP (Numba). As expected, they are significantly faster than pure Python, as Numba compiles Python code into optimized machine code, much like Cython, bypassing the slower interpreted execution of standard Python. This leads to substantial reductions in the execution times, particularly for compute-intensive tasks such as numerical computations. In any case, we have observed problems in the scalability of some benchmarks. For example, with 24 threads, the best speedups were obtained for the quad and pi applications. However, these speedups are quite limited, reaching only 2.32$\times$ and 2.27$\times$, respectively. This corresponds to a low parallel efficiency of just 9.7% and 9.5%, calculated as the ratio of speedup to the number of threads used, expressed as a percentage. fft scales only up to 4 threads, while performance degrades for higher thread counts. On the other hand, the lu and md applications do not scale at all, and their execution times increase as the number of running threads grows. Additionally, PyOMP was unable to execute the jacobi method using more than one thread.
@@ 4.2 Applications using non-numerical libraries and Python objects
Figure 12 shows the scalability of the Graph Clustering and Wordcount applications using up to 48 threads. The scalability of OMP4Py is good, especially for Wordcount, which achieves a 25.5$\times$ speedup compared to the sequential execution when using 48 threads. This behavior is very different from the performance results obtained for the numerical applications (see Figure 8). However, the structure and OpenMP directives used in both types of applications are very similar, with the only differences being related to the operations performed inside the loops. For example, when closely examining the OMP4Py codes for the $\pi$ calculation (Figure 10, top code) and the Wordcount application (Figure 13), it can be observed that the OpenMP clauses are semantically identical. In the case of a dictionary reduction, the reduction clause cannot be used, and manual implementation with a critical block is necessary. On the other hand, for numerical variables, the reduction clause can be used, and OMP4Py automatically generates the same corresponding critical block. This additional observation, together with the analysis in Section 4.1, confirms that the scalability issues detected in the $\pi$ calculation cannot be attributed to OMP4Py, as the generated code for both applications is essentially identical. Therefore, these results can only be explained by the fact that free-threaded support for Python 3.13 is still in an experimental stage, leaving plenty of room for improving performance and efficiency.
@@
Figure 13: OMP4Py user code for the Wordcount application.
@@ 5 Conclusions
In this paper, we introduce OMP4Py [^5] It is publicly available at https://github.com/citiususc/omp4py, a novel implementation of the OpenMP standard designed specifically for Python. It fully supports the API specifications of version 3.0, ensuring that we address most of the needs of OpenMP HPC applications and programmers. OMP4Py integrates OpenMP into Python by adapting its directive-based approach, allowing users to embed parallel constructs directly into their code. These transformer directives enable parallel execution by instructing OMP4Py to modify the Python code as needed.
The performance evaluation of OMP4Py reveals limited scalability when using numerical algorithms. Specifically, only 4 out of 6 benchmarks showed improvement with 4 threads, with performance typically degrading when more threads were used. However, we demonstrated that these scalability issues are due to Python’s multithreading limitations in version 3.13, rather than OMP4Py’s implementation. It is important to highlight that as these limitations in the Python interpreter are progressively resolved, the scalability constraints of OMP4Py when running numerical applications will gradually disappear.
@@
In the future, OMP4Py will be extended to support newer versions of the OpenMP standard, including versions 4.0 through 6.0. This will involve incorporating advanced features such as task dependencies, thread teams and support for accelerators. Additionally, the current implementation will be optimized by reducing mutex locks and exploring atomic operations, which are commonly used in C-based OpenMP implementations.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors just used chatGPT-3.5 in order to improve readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.